OpenStudy (anonymous):

What is the approximate distance from (-2,3) to (4,1)? 5 units 6 units 8 units 9 units

OpenStudy (amistre64):

subtract the point, square them, add them and sqrt them

OpenStudy (amistre64):

(4,1) -(-2,3) ------- (6,-2)^2 = (36 + 4) = 40 sqrt(40) = 2 sqrt(10)

OpenStudy (amistre64):

sqrt(9),sqrt(10), sqrt(16) 3, , 4 id say it close to 6 without a calculator

OpenStudy (anonymous):

thx